# Quantum-enabled optical large-baseline interferometry: applications, protocols and feasibility

### v1

@@ VI.5 Binary encoding: optical interferometry with quantum Networks [3]
The binary encoding proposed in [3] is efficient in the entanglement consumption in the parity check. Let us
@@
A encoded logical CNOT is performed between the photonic degrees of freedom and the memory qubits at each time bin
@@
This encoding keeps track of the time-of-arrival information of the photon: empty time bins leave the memories unchanged.
$\bar{\rho}_{0}\equiv\ket{\bar{0},\bar{0}}\bra{\bar{0},\bar{0}}$ | (39)
Performing the CX operation,
@@
Then the photonic degree of freedom is decoupled from the memory qubit by measuring it in the X basis. (This step is likely challenging, which is solved in the next paper). Then, applying the appropriate correction (or post-processing), the memory qubits end up being
$\rho_{\star}=(1-\epsilon)\bar{\rho}_{0}+$ $\frac{\epsilon(1+|g|)}{2}\ket{\psi^{+}_{\phi,m}}\bra{\psi^{+}_{\phi,m}}$ |
@@
where |$\ket{\psi^{\pm}_{\phi,m}}=\frac{1}{\sqrt{2}}(\ket{\bar{0},\bar{1}_{m}}\pm e^{i\phi}\ket{\bar{1}_{m},\bar{0}})$
In summary, if a photon arrives in the $(m-1)$-th time bin, then it is stored in such a way that the time bin is flipped according to |$\ket{\bar{1}_{m}}$.
This protocol uses the same set of memory for each time bin, which means that gates are applied to configure each subsequent time bin encoding. Given the short coherence time for each time bin, the gate speed needs to catch up with this operation.
VI.5.1 Error due to multiphoton contribution
Let us examine errors due to multiphoton contribution and how to alleviate them.
Starlight is thermal, and the block size is chosen such that $M\epsilon\sim 1$. If the average photon number in the block is 1, then the probability of having more than 1 photon in the block is also high. Let us examine the state post-encoding; using the notation of Ref. [3]
@@ VI.5.1 Error due to multiphoton contribution
where $p_{0}$ is the probability of receiving the vacuum, $p_{1}$ is the probability of having a single photon; and $\rho_{\text{multi}}$ is the state of the memory where more than one photon has landed. In the case of $\rho_{\text{multi}}$, the memory qubit is considered completely depolarised.
For $\gamma=1$, we have [3],
$p_{0}=\frac{1}{\epsilon+1},\qquad p_{1}=\frac{\epsilon}{(1+\epsilon)^{2}}$ (44)
If we have $\rho_{AB}^{\otimes M}$ in a block of $M$, then final decoded state in the readout qubit has coefficients given by a trinomial distribution of the input coefficients:
@@
This presents a problem, because for high-contrast $c$ needs to be as close to 1 as possible, and this is fundamental; $c$ depends on how we choose $M\epsilon$ (Fig. 7).
@@ VI.6 Quantum-assisted telescope arrays [4]
$\rho_{\nu 1}=$ $|\displaystyle\frac{1}{2}(\ket{0,1}\bra{0,1}+\ket{1,0}\bra{1,0}+g_{1}\ket{0,1}\bra{1,0}$ |
$+$ $g_{1}^{*}\ket{1,0}\bra{0,1})_{A,B,\nu 1}\otimes\ket{0,0}\bra{0,0}_{A,B,\nu 2}$
$\rho_{\nu 2}=$ |$\ket{0,0}\bra{0,0}_{A,B,\nu 1}\otimes\frac{1}{2}(\ket{0,1}\bra{0,1}+$
[…]
The notation here is:
|$\ket{\underbrace{101}_{\text{time}}\underbrace{1}_{\text{freq}},0000}$ (49)
The first 3 memory qubits encode the time bin, and the 4th encodes the frequency band.
In this particular example, integrating the stellar photon into the memory leads to the following transformation.
|$\ket{0,0}_{\nu 1}\ket{0,0}_{\nu 2}\ket{0000_{A},0000_{B}}$
$\rightarrow$ |$\ket{0,0}_{\nu 1}\ket{0,0}_{\nu 2}\ket{0000_{A},0000_{B}},$ (50)
|$\ket{1,0}_{\nu 1}\ket{0,0}_{\nu 2}\ket{0000_{A},0000_{B}}$
$\rightarrow$ |$\ket{1,0}_{\nu 1}\ket{0,0}_{\nu 2}\ket{1010_{A},0000_{B}},$ (51)
|$\ket{0,1}_{\nu 1}\ket{0,0}_{\nu 2}\ket{0000_{A},0000_{B}}$
$\rightarrow$ |$\ket{0,0}_{\nu 1}\ket{0,1}_{\nu 2}\ket{0000_{A},1010_{B}},$ (52)
|$\ket{0,0}_{\nu 1}\ket{1,0}_{\nu 2}\ket{0000_{A},0000_{B}}$
$\rightarrow$ |$\ket{0,0}_{\nu 1}\ket{1,0}_{\nu 2}\ket{1011_{A},0000_{B}},$ (53)
|$\ket{0,0}_{\nu 1}\ket{0,1}_{\nu 2}\ket{0000_{A},0000_{B}}$
$\rightarrow$ |$\ket{0,0}_{\nu 1}\ket{0,1}_{\nu 2}\ket{0000_{A},1011_{B}}.$ (54)
@@
One technical difficulty associated with using the time bin encoding is that we need to encode the memory qubits as fast as the time bins. This timescale is extremely fast compared to the current gate times for most platforms.
@@ VI.7 Incorporating quantum error correction [5]
The technically difficult part of this protocol includes:
  • The STIRAP step needs to be extremely low noise: the stellar state is unencoded at this point, and the error rate must be $\ll\epsilon$.
  • For a complete population transfer, here $t=50$, which appears long. The cavity quality needs to be sufficiently high to retain the stellar photon until the transfer is complete.
  • STIRAP schemes have very narrow bandwidth. (Rare earth: 10k Hz)
As an example of a suitable STIRAP scheme, consider a 87Rb atom trapped in an optical tweezer and coupled to a fibre Fabry-Perot cavity as recently demonstrated [74]. Adapted to the goal here of coherent single-photon transfer, we follow the prescription in [75]. Choose qubit states |$\ket{0}=\ket{5^{2}S_{1/2},F=1,m_{F}=0}$, |$\ket{1}=\ket{5^{2}S_{1/2},F=2,m_{F}=0}$, and |$\ket{e}=\ket{5^{2}P_{3/2},F=3,m_{F}=0}$, where the linewidth of the |$\ket{1}\leftrightarrow\ket{e}$ transition ($\lambda=780\,$nm) is $\gamma=2\pi\times 6\,$MHz (FWHM). We assume a cavity finesse $F\approx 2\times 10^{5}$, a waist radius ${\rm w}_{r}\approx 2\,\mu$m and a length $L\approx 40\,\mu$m resulting in a cooperativity of $C=3\lambda^{2}F/(2\pi^{3}{\rm w}_{r}^{2})\approx 1500$ (as demonstrated in [76]) with a coupling strength of $g=\sqrt{3\lambda^{2}C\gamma}/(2\pi^{2}{\rm w}_{r}^{2}L)\approx 2\pi\times 400\,$MHz and $\kappa=\pi C/LF\approx 2\pi\times 20\,$MHz (FWHM), so that $\gamma/\kappa\approx 0.3$. Over a population transfer period of $T$ one estimates a fidelity of $F=e^{-T\kappa}$ which for $T=50g^{-1}$ would be too small, $F<0.1$. Better cavities providing a factor of $10$ improvement in $g/\kappa$ would provide $F>0.75$. Note Tweezer induced dephasing rates on state |$\ket{1}$ can be treated as negligible $\gamma_{\phi}/g=0.03\times 10^{-6}$ [77].
@@ VI.8 Continuous-variable investigations
However, the advantage is very limited because it occurs for a small region of loss, and the magnitude of the advantage is also small. Practical difficulties further impede achieving any quantum advantage, limiting the merits of a CV teleportation-based strategy for stellar interferometry. Conclusion: TMSV teleportation-based strategies are practically infeasible.
@@ VI.9 Vacccum beam guide for quantum networks [6]
This paper introduces a simple and elegant solution for long-distance distribution of quantum states: a ground-based optical channel composed of a vacuum tube with precisely aligned lenses. The authors demonstrate that a VBG with realistic parameters can achieve an attenuation rate as low as $3\times 10^{-5}$ dB/km. The system avoids the complexity of quantum repeaters by requiring only passive optical elements. Key sources of loss—lens imperfections, residual gas absorption, and beam misalignment—are rigorously analysed and shown to be manageable with current technology. The system avoids the complexity of quantum repeaters by requiring only passive optical elements. Key sources of loss—lens imperfections, residual gas absorption, and beam misalignment—are rigorously analysed and shown to be manageable with current technology.
This architecture is not only promising for quantum communication but also enables transformative advances in quantum-enabled long-baseline optical interferometry (QVLBI). In particular, we see two distinct, yet complementary, directions in which the VBG infrastructure could be adapted for interferometric science.
VI.9.1 High-Efficiency Entanglement Distribution for Quantum Interferometry
In a repeater-based VLBI architecture, VBGs could serve as the physical backbone for entanglement distribution between telescope nodes. This offers several key advantages:
  • Low-loss transmission over thousands of kilometres enables high-rate distribution of entangled photons with reduced need for quantum repeaters.
  • The VBG channel can support a variety of quantum encodings (e.g., time-bin, polarisation, frequency-bin) suitable for distributed quantum metrology and interferometry.
@@ VI.9.2 Using the VBG as an interferometer arm
Alternatively, the VBG itself can be directly repurposed as an interferometric guide channel, effectively acting as one arm of a giant optical interferometer:
In this approach, starlight collected at separate telescopes is coupled into the VBG and transported to a common detection point for interference-based measurement. Since the VBG avoids the material losses and dispersion of optical fibres, it provides an ideal medium for coherent light propagation over long baselines.
However, in this configuration, one must reintroduce active variable path compensation to maintain fringe stability within a fraction of an optical wavelength. This could be achieved through delay lines and phase-locking mechanisms, possibly guided by a bright reference source or classical beacon.
This direct-interference approach would closely mimic classical optical interferometry but at scales (1000+ km) previously accessible only in the radio domain.
In addition, if implemented on the Moon, the vacuum beam guide would benefit from the natural hard vacuum of the lunar environment, eliminating the need for complex evacuated tubes. This could significantly reduce engineering overhead while enabling ultra-long baseline quantum interferometry with minimal optical loss.
@@ VII Open questions
There are open questions that remain. First, what is the best way to transfer the information in a single photon into a quantum memory? What are the trade-offs between adiabatic loading schemes (e.g. STIRAP) and faster dynamical techniques? Adiabatic loading schemes (e.g. STIRAP) offer robustness but are slow, while dynamical schemes allow faster coupling at the cost of greater sensitivity to timing and control. The optimal choice depends on memory coherence time and system timing constraints.
How many temporal modes do we actually need? Does this depend solely on $\epsilon$, or do detector and memory limitations impose stricter bounds? What should the mode shapes look like in practice? For binary encoding, the number of memory qubits scales as $\log(1/\epsilon)$, but practical constraints such as detector dead time and memory reset times also play a role.
Is temporal or wavelength multiplexing unavoidable? If so, how can we implement it without overwhelming the memory and feedforward control? Can we design architectures that minimise multiplexing while preserving sensitivity? While multiplexing boosts photon capture and entanglement rates, it increases system complexity. Minimising multiplexing while maintaining detection efficiency remains an open design question.
In summary, mode-resolved quantum VLBI imposes a unique set of challenges that bridge quantum optics, temporal mode engineering, and real-time feedback. Solving these problems will be essential for unlocking the full potential of quantum-enhanced astronomical imaging and precision geodesy.
@@ VIII Concluding remarks
This report has outlined the feasibility and potential of quantum-enabled optical Very Long Baseline Interferometry (VLBI). By integrating quantum information protocols, many long-standing challenges in optical VLBI can be mitigated.
From a technological perspective, we examined a range of architectures, including repeater-based entanglement distribution, binary-encoded quantum memories, and hybrid protocols. We believe the uses of the vacuum beam guide architecture, either as a lossless medium for entanglement distribution or as a coherent optical delay channel for direct interferometry, offer scalable paths toward kilometre-scale optical baselines.
Encouragingly, most of the individual hardware components—such as high-cooperativity optical cavities, long-lived quantum memories, and entangled photon sources—are now approaching the threshold of practical feasibility. Experimental progress in photon collection, phase stabilisation, and entanglement distribution suggests that several essential subsystems could be prototyped with near-term technologies.
The primary bottlenecks that remain are the speed and fidelity of quantum gate operations. In particular, time-bin multiplexed architectures impose stringent requirements on gate execution times and error rates that exceed current capabilities in most physical platforms. These limitations currently constrain the implementation of real-time encoding, decoding, and error correction procedures necessary for robust operation.
Nevertheless, the convergence of quantum technologies toward these benchmarks is rapid. Continued improvements in quantum control and photonic integration may soon enable the deployment of quantum-enhanced VLBI systems capable of achieving diffraction-limited resolution at optical wavelengths. Such a capability would unlock new frontiers in both astronomical imaging and precision Earth science.
@@ Acknowledgements.
We thank Max Charles, John Bartholomew, Michael Steel, Christian Schwab, John Jeffers, Lorenzo Maccone, and Chiara Macchiavello for insightful discussions. ZH is supported by an ARC DECRA Fellowship (DE230100144) “Quantum-enabled super-resolution imaging” and the EPSRC International Network in Space Quantum Technologies (EP/W027011/1). MKS is supported by an Australian Research Council Discovery Early Career Researcher Award (DE220101272). DO is supported by the EPSRC Quantum Technology Hub in Quantum Communication (EP/T001011/1), EPSRC Integrated Quantum Networks Research Hub (EP/Z533208/1), and ESA (4000147561/25/NL/FGL/ss). PK is supported by EPSRC’s Large Baseline Quantum-Enhanced Imaging Networks grant (EP/V021303/1) and Mathematical Tools for Practical Quantum Imaging Protocols grant (UKRI069). We acknowledge the traditional custodians of the Macquarie University land, the Wallumattagal clan of the Dharug nation, whose cultures and customs have nurtured and continue to nurture this land since the Dreamtime. We pay our respects to Elders past, present and future.
Appendix A Tables for reference
Wavelength $\lambda$ (nm) Frequency $\nu$ (THz) $\Delta\lambda$ (pm)
600 500 0.12
1500 200 0.75
10000 30 33
Table 4: Conversion of a frequency bandwidth of $\Delta\nu=10^{8}$ Hz to wavelength bandwidth at different central wavelengths. The coherence time is $10$ ns, as per Eq. (19).
Wavelength $\lambda$ (nm) Frequency $\nu$ (THz) $\Delta\nu$ (THz)
400 750 18.8
500 600 12.0
600 500 8.3
1000 300 3.0
1550 193.5 1.25
2000 150 0.75
Table 5: Conversion of a 10 nm wavelength bandwidth to frequency bandwidth at different central wavelengths.
Appendix B Calculating the mean photon numbers of nearby stars and exoplanets

### v2

@@ VI.5 Binary encoding: optical interferometry with quantum Networks [3]
The binary encoding proposed in [3] is efficient in memory and entanglement consumption. In this protocol, we
@@
A encoded logical CNOT (CXm) is performed between the photonic degrees of freedom and the appropriate memory qubits at each time bin
@@
This encoding keeps track of the time-of-arrival information of the photon: empty time bins leave the memories unchanged. Define $\bar{\rho}_{0}\equiv\ket{\bar{0},\bar{0}}\bra{\bar{0},\bar{0}}$ | Perform the CXm operation, we have
@@
Then the photonic degree of freedom is decoupled from the memory qubit by measuring it in the X basis. This step is likely challenging, which can be solved in the next protocol. Then, applying the appropriate correction, the memory qubits end up in the state
$\rho_{AB}\rightarrow(1-\epsilon)\bar{\rho}_{0}+$ $\frac{\epsilon(1+|g|)}{2}\ket{\psi^{+}_{\phi,m}}\bra{\psi^{+}_{\phi,m}}$ |
@@
where |$\ket{\psi^{\pm}_{\phi,m}}=\frac{1}{\sqrt{2}}(\ket{\bar{0},\bar{1}_{m}}\pm e^{i\phi}\ket{\bar{1}_{m},\bar{0}})$.
In summary, if a photon arrives in the $(m-1)$-th time bin, then it is stored in such a way that the time bin is prepared according to |$\ket{\bar{1}_{m}}$.
This protocol uses the same set of memory qubits for all time bins, which require gates to be applied in between each subsequent time bin encoding: the implication is that the gate speed needs to catch up with the time bin size.
Errors due to multiphoton contribution
Let us examine errors due to multiphoton contribution in this protocol and how to alleviate them.
Starlight is thermal, if we choose the time bin block size such that $M\epsilon\sim 1$, then the mean photon number in the block is 1; if so, the probability of having more than 1 photon in the block is also non-negligible. Examine the state post-encoding:
@@
where $p_{0}$ is the probability of receiving the vacuum, $p_{1}$ is the probability of having a single photon, and $\rho_{\text{multi}}$ is the state of the memory where more than one photon has landed. In the case of $\rho_{\text{multi}}$, the memory qubit is considered completely depolarised.
For $\gamma=1$, $p_{0}={1}/(\epsilon+1)$, $p_{1}={\epsilon}/{(1+\epsilon)^{2}}$ [3]. If we have $\rho_{AB}^{\otimes M}$ in a block of $M$, then final decoded state in the readout qubit has coefficients given by a trinomial distribution of the input coefficients:
@@
This presents a problem, because for high-contrast $c$ needs to be as close to 1 as possible, and this is fundamental; $c$ depends on how we choose $M\epsilon$ (Fig. 7). If we frequency multiplex the signal such that it is spread over more memory qubits, the probability of infidelity decreases.
Entanglement consumption rate
@@ VI.6 Quantum-assisted telescope arrays [4]
The notation here for the 2nd frequency band is:
|$\underbrace{\ket{0_{A_{5}},0_{B_{5}}}}_{\text{freq 1}}\underbrace{\ket{1_{A_{5}},0_{B_{5}}}}_{\text{freq 2}}\underbrace{\ket{1011_{A},0000_{B}}}_{\text{memories}}$
|$\underbrace{\ket{0_{A_{5}},0_{B_{5}}}}_{\text{freq 1}}\underbrace{\ket{0_{A_{5}},1_{B_{5}}}}_{\text{freq 2}}\underbrace{\ket{0000_{A},1011_{B}}}_{\text{memories}}.$ (45)
The first 3 memory qubits encode the time bin, and the 4th encodes the frequency band:
|$\ket{\underbrace{101}_{\text{time}}\underbrace{1}_{\text{freq}}{}_{A},0000_{B}}$ (46)
@@ VI.7 Incorporating quantum error correction [5]
The technically difficult part of this protocol may include: 1) The STIRAP step needs to be extremely low noise: the stellar state is unencoded at this point, and the error rate must be $\ll\epsilon$. 2) For a complete population transfer, here $t=50$, which appears long. The cavity quality needs to be sufficiently high to retain the stellar photon until the transfer is complete. e) STIRAP schemes tend to have very narrow bandwidth, in rare Earth ions the typical bandwidth is around 10k Hz
As an example of a suitable STIRAP scheme, consider a 87Rb atom trapped in an optical tweezer and coupled to a fibre Fabry-Perot cavity as recently demonstrated [78]. Adapted to the goal here of coherent single-photon transfer, we follow the prescription in [79]. Choose qubit states |$\ket{0}=\ket{5^{2}S_{1/2},F=1,m_{F}=0}$, |$\ket{1}=\ket{5^{2}S_{1/2},F=2,m_{F}=0}$, and |$\ket{e}=\ket{5^{2}P_{3/2},F=3,m_{F}=0}$, where the linewidth of the |$\ket{1}\leftrightarrow\ket{e}$ transition ($\lambda=780\,$nm) is $\gamma=2\pi\times 6\,$MHz (FWHM). We assume a cavity finesse $F\approx 2\times 10^{5}$, a waist radius ${\rm w}_{r}\approx 2\,\mu$m and a length $L\approx 40\,\mu$m resulting in a cooperativity of $C=3\lambda^{2}F/(2\pi^{3}{\rm w}_{r}^{2})\approx 1500$ (as demonstrated in [80]) with a coupling strength of $g=\sqrt{3\lambda^{2}C\gamma}/(2\pi^{2}{\rm w}_{r}^{2}L)\approx 2\pi\times 400\,$MHz and $\kappa=\pi C/LF\approx 2\pi\times 20\,$MHz (FWHM), so that $\gamma/\kappa\approx 0.3$. Tweezer-induced dephasing rates on state |$\ket{1}$ can be treated as negligible $\gamma_{\phi}/g=0.03\times 10^{-6}$ [81], so the dominant source of error would be cavity decay. Over a population transfer period of $T$, with on average $1/2$ photon occupation in the cavity, one estimates a fidelity of $F=e^{-T\kappa/2}$ which for $T=50g^{-1}$ would imply $F\approx 0.29$. Better cavities providing a factor of $5$ improvement in $g/\kappa$ would enable $F>0.77$, which could be further improved by post-selecting on null results from single photon detectors monitoring leaked photons [82].
@@ VI.8 Continuous-variable investigations
However, the advantage is very limited because it occurs for a small region of loss, and the magnitude of the advantage is also small. Practical difficulties further impede achieving any quantum advantage, limiting the merits of a CV teleportation-based strategy for stellar interferometry. Consistent with Ref. [69], we conclude that TMSV teleportation-based strategies are practically infeasible if there is no quantum repeater available.
@@ VI.9 Vacccum beam guide for quantum networks [6]
This paper introduces a simple and elegant solution for long-distance distribution of quantum states: a ground-based optical channel composed of a vacuum tube with precisely aligned lenses. The authors demonstrate that a VBG with realistic parameters can achieve an attenuation rate as low as $3\times 10^{-5}$ dB/km. The system avoids the complexity of quantum repeaters by requiring only passive optical elements. Key sources of loss—lens imperfections, residual air absorption, and beam misalignment—are analysed and shown to be manageable with current technology.
This architecture is not only promising for quantum communication but also enables advances in quantum-enabled O-VLBI. In particular, we see two distinct, yet complementary, directions in which the VBG infrastructure could be adapted:
VI.9.1 High-Efficiency Entanglement Distributor
In a repeater-based VLBI architecture, VBGs could serve as the physical backbone for entanglement distribution between telescope nodes. This offers several advantages. Low-loss transmission over thousands of kilometres enables high-rate distribution of entangled photons with reduced need for quantum repeaters. The VBG channel can support a variety of quantum encodings (e.g., time-bin, polarisation, frequency-bin) suitable for distributed quantum metrology and interferometry.
@@ VI.9.2 Using the VBG as an interferometer arm
Alternatively, the VBG itself can be directly repurposed as an interferometric guide channel, effectively acting as one arm of a large optical interferometer. In this approach, starlight collected at separate telescopes is coupled into the VBG and transported to a common detection point for interference-based measurement. Since the VBG avoids the material losses and dispersion of optical fibres, it provides an ideal medium for coherent light propagation over long baselines. However, in this configuration, one must reintroduce active variable path compensation to maintain fringe stability within a fraction of an optical wavelength. This could be achieved through delay lines and phase-locking mechanisms, possibly guided by a bright reference source or classical beacon. In addition, if implemented on the Moon, the vacuum beam guide would benefit from the natural hard vacuum of the lunar environment, eliminating the need for complex evacuated tubes. This could significantly reduce engineering overhead with minimal optical loss.
@@ VII Open questions
There are open questions that remain. First, what is the best way to transfer the information in a single photon into a quantum memory? What are the trade-offs between adiabatic loading schemes (e.g. STIRAP) and faster dynamical techniques? Adiabatic loading schemes offer robustness but are slow, while dynamical schemes allow faster coupling at the cost of having high noise to timing and control. The optimal choice depends on memory coherence time and system timing constraints.
How many temporal modes do we actually need? Does this depend solely on $\epsilon$, or do detector and memory limitations impose stricter bounds? What should the mode shapes look like in practice? For binary encoding, the number of memory qubits scales as $\log(1/\epsilon)$, but practical constraints such as detector dead time and memory reset times also play a role. How can we most efficiently implement multiplexing without overwhelming the memory and feedforward control?
The atmosphere fluctuates on the timescale of milliseconds. What is the best way of correcting for the phase variations? What requirements are needed for a potential artificial guiding star? Can we collect enough photons by using broadband light to avoid using a guiding star?
How might established techniques such as closure-phase observables be adapted within quantum-enhanced architectures? More broadly, can quantum estimation theory enable new imaging schemes that go beyond pairwise visibility measurements and the standard van Cittert–Zernike framework?
In the absence of full quantum repeaters, what level of VLBI enhancement is realistically achievable using repeater-less quantum links? Can the distribution of additional ground photons, particularly with more robust temporal or mode entanglement, meaningfully improve signal-to-noise ratio? Furthermore, how might such protocols connect with or inform developments in all-optical quantum repeater architectures?
Beyond classical Fisher information and maximum likelihood methods, could quantum-enhanced or learning-assisted inference techniques play a role in astronomical imaging? Recent work at the interface of quantum estimation and learning theory? E.g., Ref. [86] suggests a potentially fruitful direction for developing robust, data-efficient reconstruction techniques tailored to interferometric observations.
@@ VIII Concluding remarks
This report has outlined the feasibility and potential of quantum-enabled optical VLBI. By integrating quantum information protocols, many long-standing challenges in optical VLBI can be mitigated.
From a technological perspective, we examined a range of architectures. Encouragingly, most of the individual hardware components—such as high-cooperativity optical cavities, long-lived quantum memories, and entangled photon sources—are now approaching the threshold of practical feasibility. Experimental progress in photon collection, phase stabilisation, and entanglement distribution suggests that several essential subsystems could be prototyped with near-term technologies. The primary bottlenecks that remain are the speed and fidelity of quantum gate operations. In particular, time-bin multiplexed architectures impose stringent requirements on gate execution times and error rates that exceed current capabilities in most physical platforms.
Nevertheless, the development of quantum technologies toward these benchmarks is rapid. Continued improvements in quantum control and photonic integration may soon enable the deployment of quantum-enhanced VLBI systems capable of achieving diffraction-limited resolution at optical wavelengths. Such a capability would unlock new frontiers in both astronomical imaging and precision Earth science.
@@ Acknowledgements.
We thank Max Charles, John Bartholomew, Michael Steel, Christian Schwab, John Jeffers, Lorenzo Maccone, and Chiara Macchiavello, Yujie Zhang, Yunkai Wang and Virginia Lorenz for insightful discussions. ZH is supported by an ARC DECRA Fellowship (DE230100144) “Quantum-enabled super-resolution imaging” and the EPSRC International Network in Space Quantum Technologies (EP/W027011/1). MKS is supported by an Australian Research Council Discovery Early Career Researcher Award (DE220101272). DO is supported by the EPSRC Quantum Technology Hub in Quantum Communication (EP/T001011/1), EPSRC Integrated Quantum Networks Research Hub (EP/Z533208/1), and ESA (4000147561/25/NL/FGL/ss). PK is supported by EPSRC’s Large Baseline Quantum-Enhanced Imaging Networks grant (EP/V021303/1) and Mathematical Tools for Practical Quantum Imaging Protocols grant (UKRI069). We acknowledge the traditional custodians of the Macquarie University land, the Wallumattagal clan of the Dharug nation, whose cultures and customs have nurtured and continue to nurture this land since the Dreamtime. We pay our respects to Elders past, present and future.
Appendix A Calculating the mean photon numbers of nearby stars and exoplanets